OpenStudy (anonymous):

Simplify: log base 3 (1/3)

OpenStudy (anonymous):

my head hurts.... sorry i not helpful

jimthompson5910 (jim_thompson5910):

\[\large \log_{3}\left(\frac{1}{3}\right)\] \[\large \frac{\log_{10}\left(\frac{1}{3}\right)}{\log_{10}\left(3\right)}\] \[\large \frac{\log_{10}\left(3^{-1}\right)}{\log_{10}\left(3^1\right)}\] \[\large \frac{-1*\log_{10}\left(3\right)}{1*\log_{10}\left(3\right)}\] \[\large \frac{-1}{1}\] \[\large -1\] So \[\large \log_{3}\left(\frac{1}{3}\right)=-1\]

OpenStudy (anonymous):

wow. ^ he is SMART!!!! (: u on fb smarty pants??? ;)

hero (hero):

It wasn't that hard

jimthompson5910 (jim_thompson5910):

Quick shortcut is to type log(1/3)/log(3) into your calculator and you'll get the same answer.

OpenStudy (anonymous):

lol yeah he really is! He's helped me so much. Thanks guys! Well it's not the easiest thing for me, I barely remember how to do this stuff so

OpenStudy (anonymous):

Can I do that with all logs?

jimthompson5910 (jim_thompson5910):

Yes, this trick works with any log.

OpenStudy (anonymous):

^ hey jim_thompson5910 can u help on my problem? i just asked it. merci beacoup if you can look at it! (: